OpenStudy (anonymous):

if 14% of the group voted to change its bylaws and 16% voted ageinst it , how many people did not vote ? the group has 460 members

OpenStudy (ivettef365):

that is a total of 30% that voted, so what is the difference

OpenStudy (anonymous):

0.16% ??

OpenStudy (ivettef365):

a whole thing is 100%, so if 30% voted, how many did not vote

OpenStudy (texaschic101):

14% voted to change -- (.14)(460) = 64.4 16% voted to not change it -- (.16)(460) = 73.6 total : 30% = 64.4 + 73.6 = 138 Leaving this many people not voting 460 - 138 = 322 Or, like ivette said.....30% voted....leaving 100% - 30% = 70% who didn't vote. And 70% of 460 = (.70)(460) = 322 Either way you do it, you will get the same answer....322 people didn't vote.

OpenStudy (anonymous):

(14% + 16 % )/100 * 460 - 460 = 322

OpenStudy (anonymous):

3w * w = 48 write the value of w in simplist form .. help ?